7x+6=6
7X=6-6
7X=0
X=?
Is the answer 7 or 0?

Answer by Alan3354(69443)

7x+6=6
7X=6-6
7X=0
X=?
Is the answer 7 or 0?
===========
7x = 0
Divide by 7
x = 0/7
x = 0